# Large Language Model Can Continue Evolving From Mistakes

### v5

@@ Abstract
As world knowledge evolves and new task schemas emerge, Continual Learning (CL) is crucial for keeping Large Language Models (LLMs) up-to-date and addressing their shortcomings. LLMs typically require continual instruction tuning (CIT) and continual pre-training (CPT) to adapt to new tasks and acquire essential knowledge. However, collecting sufficient CPT data while addressing knowledge gaps remains challenging, as does optimizing the efficiency of utilizing this data. Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To enhance data utilization and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT data. Experiments demonstrate that CEM significantly enhances model performance and continual evolution. The code and dataset are available in the GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
Large Language Model Can Continue Evolving From Mistakes
@@ 1 Introduction
With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs. Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al. (2023). Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial. Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al. (2024).
@@
To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al. (2022); Jin et al. (2022). However, these methods lack focus and efficiency and fail to dynamically update model deficiencies. Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al. (2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al. (2024) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data. Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation. To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’. We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies. Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes for directly address these deficiencies. To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data. In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution. The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 17.00% accuracy enhancement.
@@ 2 Related Work
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3 Methods
3.1 Task Formulation
Question Answering (QA) tasks are commonly used to evaluate LLM’s domain knowledge. These tasks typically involve a series of single-choice questions with multiple options. If an LLM correctly answers these questions, it indicates a strong grasp of the relevant knowledge. Conversely, incorrect answers reveal deficiencies in understanding.
Given a question-answer example $(qi,ai)\in\mathcal{Q}$, a test set for question answering $\mathcal{Q}$, and a reasoning model $\mathcal{M}_{0}$.
First, the reasoning model $\mathcal{M}_{0}$ is utilized to answer all questions within the test set $\mathcal{Q}$. Subsequently, questions $q_{i}$ that receive incorrect responses are collected,
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
where $\widetilde{r_{i}}$ denotes the response that includes incorrect answers to the question $q_{i}$, and $Ans(\cdot)$ is utilized to extract the final answer from the response.
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$ as CPT data,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\ldots\},$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
@@ 3.3 Supplemental Training Data
Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
@@
Below are the three types of CIT data:
(2) Normative Corpus($C_{\underline{N}}$):
To enhance the LLM’s understanding of task schemas and formats, task-relevant instruction fine-tuning is commonly applied. For QA tasks, we sample questions from the training set and format them as shown in Table 3 to build the Normative Corpus, enabling the LLM to answer multiple-choice questions in a standardized manner. Before supplemental training, we fine-tune the LLM using the Normative Corpus (i.e., pre-finetuning). A small portion of the Normative Corpus is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample questions from the training set and format them as shown in Table 4 to build the Extractive Corpus. We employ Qwen2-72B-Instruct for entailment checking to filter the corpus. A minimal amount of Extractive Corpus data yields effective results, ensuring manageable inference costs. Details are provided in the Appendix E.
(4) Review Corpus($C_{\underline{R}}$):
@@
To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round. We process some of these responses in the Normative Corpus format and the rest in the Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus is described as follows:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^8] https://huggingface.co/datasets/Aeala/ShareGPT_
@@
as the General Corpus to preserve the model’s general capabilities on other tasks and instruction-following ability.
Supplemental Training Set Construction Strategy. We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data. Specifically, we construct three types of supplemental training sets by combining five types of corpora. The following sections detail these construction strategies.
(1) Plain Strategy:
$D_{\underline{P}}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges these corpora, and $D_{P}$ represents the supplemental training set constructed using the Plain Strategy.
(2) Extractive Strategy:
$D_{\underline{E}}=Concat(C_{S},C_{N},C_{G},{\color[rgb]{1,0,0}C_{E}}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ represents the supplemental training set constructed using the Extractive Strategy.
(3) Review Strategy:
$D_{\underline{R}}=Concat(C_{S},C_{N},C_{G},C_{E},{\color[rgb]{1,0,0}C_{R}}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ represents the supplemental training set constructed using the Review Strategy.
@@ 4.1 Training Setup
We train the models on 2 A800 GPUs using ZeRO (Rajbhandari et al., 2020) stage 1 from DeepSpeed (Rasley et al., 2020). We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 16 with a maximum sequence length of 2,048 and 1 training epoch.
@@ 4.2 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a).
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities. Appendix G shows the details of the data division and usage.
@@ 4.3 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs include CuteGPT-13B [^9] https://huggingface.co/XuYipei/kw-cutegpt-13b-base Liang et al. (2023), ChatGLM-6B [^10] https://huggingface.co/THUDM/chatglm3-6b Du et al. (2022), and Qwen-7B-Chat [^11] https://huggingface.co/Qwen/Qwen-7B-Chat Bai et al. (2023). These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
@@ 4.4 Metrics
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of taks $i$ after applying CEM on task $k$.
@@ 5 Experiments
We construct supplemental training sets and control groups for experiments as outlined in Section 3.3. To differentiate models trained on various supplemental sets, we employ the following naming conventions:
INIT represents the model pre-finetuned solely with the Normative Corpus.
CEM-P: ‘P’ for Plain, indicates the model trained using the supplemental training set constructed by the Plain Strategy.
@@
Suffixes ‘Wiki’, ‘Bing’, and ‘Mix’ indicate the data sources of the Supplemental Corpus, with ‘Mix’ signifying a combined set from Bing and Wikipedia. For instance, ‘EWiki’ denotes the model trained using the Extractive Strategy with a Supplemental Corpus from Wikipedia.
@@ 5.1 Main Results
We use the Plain Strategy to construct supplemental training set, using approximately 4,000 CPT data (based on the initial number of mistakes) and 4,000 CIT data for single-round CEM. The main results are shown in Figure 3. Our observations are summarized as follows:
Figure 3: This figure shows the accuracy of different models using the CEM-P method on the Xiezhi and CMMLU tasks.
Firstly, after supplemental training, a marked improvement in accuracy is noted across all models for both tasks. This suggests that the CEM method effectively addresses the model’s knowledge deficiencies, despite the limited training set and readily accessible supplemental data. For example, the ChatGLM-6B exhibits an increase of up to 16.08% in accuracy on the CMMLU task.
Secondly, the effectiveness of CEM varies with the innate abilities and parameter scales of LLMs. For example, on the Xiezhi task, ChatGLM-6B, with the fewest parameters, improves by up to 13.17%. In contrast, CuteGPT-13B, with the highest initial accuracy and most parameters, shows a smaller improvement of up to 3.63%.
Construction Strategy Analysis. To investigate the effectiveness of each component in the supplemental training set, we conducted experiments on CuteGPT-13B with various dataset construction strategies. We set the volume of both Extractive Corpus and Review Corpus to 1,000. The results are shown in Figure 4. Below is a summary of our observations:

### v6

# CEM: A Data-Efficient Method for Large Language Models to Continue Evolving From Mistakes

@@ Abstract
As world knowledge advances and new task schemas emerge, Continual Learning (CL) becomes essential for keeping Large Language Models (LLMs) current and addressing their shortcomings. This process typically involves continual instruction tuning (CIT) and continual pre-training (CPT) to enable these models to adapt to novel tasks and acquire critical knowledge. However, collecting sufficient CPT data and efficiently bridging knowledge gaps remain significant challenges. Inspired by the ‘summarizing mistakes’ strategy, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient approach aiming to collect CPT data and continually improve LLMs’ performance through iterative evaluation and supplementation with mistake-relevant knowledge. To further optimize data usage and mitigate forgetting, we introduce a novel training paradigm that combines CIT and CPT. Experiments show that CEM substantially enhances multiple models’ performance on both in-domain and out-of-domain QA tasks, achieving gains of up to 29.63%. Code and datasets are available on GitHub  [^1] https://anonymous.4open.science/r/cem-BB25.
CEM: A Data-Efficient Method for Large Language Models to Continue Evolving From Mistakes
@@ 1 Introduction
With the exponential growth of data and model sizes, Large Language Models (LLMs) have demonstrated superior performance across numerous downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the continual emergence of new knowledge and evolving task requirements necessitate ongoing updates and task-specific adaptations for LLMs. Without these, models risk inducing hallucinations due to knowledge boundaries and may cause task misalignment Huang et al. (2023a). Additionally, addressing and correcting the shortcomings and errors exposed during practical use is crucial. Therefore, Continual Instruction Tuning (CIT) and Continual Pre-training (CPT) are proposed as the primary methods of Continual Learning (CL) to align LLMs with evolving knowledge and tasks, and to improve their shortcomings Wu et al. (2024).
@@
To address these issues, the most common data collection methods involve using the latest Wikipedia snapshots or new academic papers in specialized fields Cossu et al. (2022); Jin et al. (2022). However, these methods lack focus and efficiency and fail to dynamically update model deficiencies. Additionally, recent work has explored using highly relevant QA pairs for CIT before CPT Jiang et al. (2024) or converting raw CPT corpus into reading comprehension instructions (i.e., CIT data) Cheng et al. (2024b) to mitigate degradation and enhance the efficiency of knowledge absorption from CPT data. Nevertheless, reliance on oracle LLMs and the complexity of mining patterns limit their practicality.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide (1) a practical pipeline for collecting CPT data based on model mistakes and (2) a novel training set construction paradigm to efficiently utilize these data, enabling continual model improvement while mitigating degradation. To enhance the targetedness and efficiency of data collection, we draw inspiration from the human learning skill of ‘summarizing mistakes’. We argue that a model’s mistakes in ‘exams’ reflect its inherent knowledge deficiencies. Therefore, we collect corpora from the internet that contain background knowledge related to these mistakes to directly address these deficiencies. To efficiently utilize these CPT data and mitigate forgetting, we design a novel training set construction paradigm that integrates parallel CIT and CPT data. In addition, by integrating CEM with methods to mitigate catastrophic forgetting (e.g., Random Replay) Li et al. (2022), we enable iterative and continual model evolution. The CEM method consistently demonstrates significant improvements across multiple models in extensive experiments, achieving up to a 29.63% accuracy enhancement.
@@ 2 Related Work
Domain-incremental CIT aims to continually fine-tune LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024b) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3.2 Supplemental Training Data
Definition of Data Type. The construction of the supplemental training set involves the following five types of data.
@@
Below are the four types of CIT data:
(2) Normative Instruction($I_{\underline{N}}$):
To enhance the LLM’s understanding of task schemas and formats, task-relevant instruction fine-tuning is commonly applied. For QA tasks, we sample questions from the training set and format them as shown in Table 4 to build the Normative Instruction, enabling the LLM to answer multiple-choice questions in a standardized manner. Before supplemental training, we fine-tune the LLM using the Normative Instruction (i.e., pre-finetuning). A small portion of the Normative Instruction is incorporated during supplemental training to maintain the model’s ability to provide normative answers.
(3) Extractive Instruction($I_{\underline{E}}$): Inspired by the human learning process, where despite studying from the same textbooks, students who excel at extracting knowledge achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To enhance the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we restructured a subset of training set questions into an information-extraction question format, as illustrated in Table 5. We employ Qwen2-72B-Instruct for entailment checking to filter the instructions. A minimal amount of Extractive Instruction yields effective results, ensuring manageable inference costs. Details are provided in the Appendix E.
(4) Review Instruction($I_{\underline{R}}$):
@@
To reinforce the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly by the LLM in the previous round. We process some of these responses in the Normative Instruction format and the rest in the Extractive Instruction format to build the Review Instruction.
The composition of the Review Instruction is defined as:
$I_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$ represents the proportion of responses in the Normative Instruction format, and $r$ denotes the correct responses from the previous round. $Normative(\cdot)$ and $Extractive(\cdot)$ convert responses into the Normative and Extractive Instruction formats, respectively. In subsequent experiments, we examine the effect of varying $\alpha$.
(5) General Instruction($I_{\underline{G}}$): We replay previous IFT data to mitigate catastrophic forgetting when training LLMs on new data He et al. (2021). When previous IFT data is unclear, we use the IFT data from the ShareGPT dataset [^8] https://huggingface.co/datasets/Aeala/ShareGPT_
@@
as the General Instruction to preserve the model’s general capabilities on other tasks and instruction-following ability.
Supplemental Training Set Construction Strategy. We propose a novel CL dataset construction paradigm that parallelizes CIT and CPT data. Specifically, we develop three supplemental training set construction strategies—Plain, Extractive, and Review—by concatenating some or all of five types of data, as shown in Step 5 of Figure 2. These strategies are named to reflect the specific data components incorporated.
@@ 4.1 Training Setup
We train the models on 2 A800 GPUs using ZeRO (Rajbhandari et al., 2020) stage 1 from DeepSpeed (Rasley et al., 2020). We adopt AdamW (Kingma and Ba, 2015) as the optimizer and set the batch size to 32 with a maximum sequence length of 2,048 and 1 training epoch.
@@ 4.2 Datasets
We conduct experiments on two question-answering datasets, Xiezhi Gu et al. (2023) and CMMLU Li et al. (2024a). To evaluate CEM’s out-of-domain generalization, we apply CEM on the in-domain dataset Xiezhi and use CMMLU and C-Eval Huang et al. (2023b) as out-of-domain benchmarks.
To investigate the impact of catastrophic forgetting, we also use the HotpotQA Yang et al. (2018) and GSM8K Cobbe et al. (2021) datasets to test the model’s decline in reasoning and mathematical abilities.
@@ 4.3 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, particularly for smaller parameter-scale models, we select open-source LLMs, including Qwen1.5-7B-Chat [^9] https://huggingface.co/Qwen/Qwen1.5-7B-Chat Bai et al. (2023), Llama3-8B-Instruct [^10] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024), and CuteGPT-13B-ift [^11] https://huggingface.co/XuYipei/kw-cutegpt-13b-ift Liang et al. (2023). Each model has been fine-tuned on the IFT data to ensure it effectively follows instructions.
@@ 4.4 Metrics
AFR: The Average Forgetting Rate Wang et al. (2024) measures the average relative degradation brought by CEM on tasks, excluding the task using CEM, compared to their initial performance. It is calculated by $AFR=\frac{1}{N-1}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{N}\left(\frac{R_{i}^{0}-R_{i}^{*}}{R_{i}^{0}}\right)$, where $R^{0}_{i}$ is the initial accuracy of task $i$ after the initial fine-tuning, and $R_{i}^{*}$ is the accuracy of tasks $i$ after applying CEM on task $k$.
@@ 5 Experiments
We construct supplemental training sets and control groups for experiments as outlined in Section 3.2. To distinguish models trained on different supplemental sets, we use the following naming conventions:
INIT represents the model pre-finetuned solely with the Normative Instruction, equipping it with the ability to generate standardized responses to multiple-choice questions. All subsequent supplemental training is conducted based on this model.
CEM-P: P for Plain, referring to the model trained using the supplemental set constructed with the Plain Strategy.
@@
We also consider several baselines for comparison with our algorithm as follows:
TempWiki: TemporalWiki uses Wikipedia snapshots as CPT data Jang et al. (2022). We construct the supplemental training set using corpus from TemporalWiki [^12] https://huggingface.co/datasets/seonghyeonye/TemporalWiki, broadly associated with the evaluation set rather than focusing solely on error cases.
AdaptLLM: AdaptLLM continues to train the LLMs on the reading comprehension texts constructed based on the raw corpora, mixed with general instructionsCheng et al. (2024b).
InstructPT: Instruction Pre-Training utilizes an instruction synthesizer [^13] https://huggingface.co/instruction-pretrain to convert raw texts from pre-training corpora into instruction-response pairs for continual tuning Cheng et al. (2024a).
@@ 5.1 Main Results
We apply the CEM method on the Xiezhi task, collecting approximately 60,000 supplemental corpus samples sourced from Wikipedia. From this corpus, we downsample 25,000 samples as CPT data, while the remaining four types of CIT data are each set to 2,000 samples. To ensure fair comparisons, we add 2,000 Normative Instructions to each of the other baselines, keeping the quantity of raw corpus data consistent. Table 1 presents the experimental results, with CMMLU and C-Eval serving as out-of-domain (OOD) benchmarks to evaluate the model’s performance on unseen QA tasks. Our observations are summarized as follows:
Benchmark INIT Ours Others
Acc(%) Prompt CEM-P CEM-E CEM-R1 TempWiki AdaptLLM InstructPT
Qwen1.5-7B-Chat
Xiezhi 42.41 51.52 52.12 56.87 30.15 35.14 49.05
CMMLU 36.08 36.30 37.44 40.34 24.22 21.12 38.50
C-Eval 35.00 37.30 40.34 40.90 27.10 16.50 36.80
Llama3-8B-Instruct
Xiezhi 25.09 35.98 40.52 54.72 32.24 41.63 42.46
CMMLU 23.00 24.92 29.76 33.92 23.40 27.62 31.30
C-Eval 26.00 26.20 30.20 31.40 22.90 25.30 32.50
CuteGPT-13B-ift
Xiezhi 45.70 56.82 57.03 58.67 50.80 45.67 54.04
CMMLU 32.56 37.96 37.42 38.26 33.44 32.32 35.22
C-Eval 31.50 35.00 35.20 36.10 31.60 28.50 33.00
Table 1: The main experimental results of our methods and baseline approaches across various tasks are presented. Experiments are conducted using three different LLMs: Qwen1.5-7B-Chat, Llama3-8B-Instruct and CuteGPT-13B-ift. The top two performances are highlighted in red bold and black bold, respectively.
Firstly, CEM leads to significant accuracy improvements across all models and tasks. All variations of CEM substantially enhance model performance in both the in-domain Xiezhi task and OOD tasks. For instance, the Llama3-8B-Instruct model achieves accuracy gains of up to 29.63% on the Xiezhi task, 10.92% on the CMMLU task, and 5.40% on the C-Eval task when using CEM-R. This highlights the effectiveness of the CEM method in addressing the model’s knowledge gaps and its capacity to generalize this knowledge to unseen QA tasks.
Secondly, CEM demonstrates superior stability compared to other baselines. Both TempWiki and AdaptLLM occasionally result in performance degradation for LLMs. For instance, AdaptLLM reduces the accuracy of the Qwen1.5-7B-Chat model by 7.27% on the Xiezhi task, 14.96% on the CMMLU task, and 18.50% on the C-Eval task. In contrast, all variations of CEM consistently deliver positive performance improvements.
Thirdly, Extractive Instruction and Review Instruction further enhance performance. A clear trend of incremental performance improvement is observed from CEM-P to CEM-E and finally to CEM-R. The ablative properties introduced during dataset construction suggest that these two types of CIT data play a crucial role in driving the observed improvements.
5.2 Data Source Analysis
We analyze the impact of the Supplementary Corpus from different sources—Wikipedia, Bing, and their concatenation—on CEM’s performance. To further validate CEM’s generalizability, CMMLU is included as an in-domain task scenario with data proportions consistent with the main experiment.
Figure 3: This figure presents the accuracy of models using the CEM-P method with different data sources on Xiezhi and CMMLU tasks. The suffixes Wiki, Bing, and Mix indicate the sources of the Supplementary Corpus, with Mix representing using double the samples sourced from the combined dataset of Wikipedia and Bing.
We find that the CEM method shows robust compatibility across different data sources, consistently improving model performance. However, integrating multiple sources does not necessarily outperform using a single source. As shown in Figure 3, no significant differences in CEM effectiveness emerge when employing Wikipedia, Bing, or their combination. Although the Mix dataset is larger, it often yields lower accuracy than Wiki or Bing, potentially due to data redundancies. Moreover, an excessively large supplemental corpus can dilute the influence of CIT data, reducing its capacity to mitigate catastrophic forgetting, support knowledge extraction, and prevent overfitting.
5.3 Mechanistic Analysis of Extractive and Review Instruction Components
To clarify the underlying mechanisms by which Extractive Instruction and Review Instruction affect model performance, we conducted experiments on CuteGPT-13B-ift and examined the R2W and W2R metrics. Additionally, we explored the effects of varying $\alpha$ (with values $[0,0.5,1]$) on CEM-R to determine whether more diverse and complex Review Instructions yield improved results. The corresponding results are illustrated in Figure 4 and Figure 5.
We summarize the key observations and insights derived from these analyses:
(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance. As shown in Figure 5, the Extractive Instruction enhances W2R performance: compared to CEM-P, CEM-E achieves improvements of 1.45%, 3.22%, and 1.26% across the three models, indicating more effective error correction. In addition, the Review Instruction significantly reduces the R2W metric, reflecting a decreased tendency for the model to forget previously acquired correct knowledge.
While the performance benefits of the Review Instruction are readily understandable, the exact mechanism by which the Extractive Instruction improves the model’s ability to integrate crucial knowledge from lengthy passages warrants further discussion.
We propose the following explanation: The Extractive Instruction provides the model with both the questions and supplementary materials containing relevant information. Questions and answer options typically present key points in a concise manner, whereas the supplementary materials often contain a dense network of factual statements Jiang et al. (2024). Thus, the questions act as a structured induction of the supplementary materials, explicitly linking complex information to the targeted induction (i.e., the questions and options). Serving as cues and guides, the questions and options help the model retrieve pertinent information from the corpus while filtering out redundant or irrelevant details. This process enhances the model’s ability to efficiently extract and utilize essential information.
(1) CEM-E and CEM-R outperform their ablation groups with General Instruction, demonstrating the effectiveness and positive impact of Extractive and Review Instruction on model performance. Figure 5 shows that Extractive Instruction improves W2R, with CEM-E achieving gains of 1.45%, 3.22%, and 1.26% across the three models compared to CEM-P, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Instruction significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
(2) The proportion factor $\boldsymbol{\alpha}$ has minimal impact on the effectiveness of Review Instruction, as R0, R0.5, and R1 exhibit similar effects on LLM performance. This indicates that reviewing knowledge, regardless of format, effectively consolidates it.